Use the following information to answer questions 9 and 10: Custom Banners manufactures advertising banners Custom allcates variable manufacturing overhead costs based on machine hours Custom budgets 0.4 machine hours per banner and llocates variable overhead at a budgeted rate of $18.00 per machine hour. Customs actual results or the curentyear were as follows: manufactured 32,250 banners,used 12000 machine hours, and reported a favorable variable manufacturing overhead rate variance of $15,000.
media%2Fc4b%2Fc4befbfb-c411-451f-a51d-60
Answer #1

Variable overhead rate variance = (Standard rate-actual rate)actual hours

15000 = (18*12000-12000X)

12000X = 216000-15000

X(actual rate) = 16.75 per hour

Actual overhead = 16.75*12000 = 201000

So answer is a) $201000